Research has shown that 55% of new Small Medium Enterprises (SMEs) are started by graduates while 45% are started by non-graduates. It is also known that 70% of SMEs started by graduates are successful i.e. they survive beyond 3 years, while only 10% of those started by non-graduates are successful.

Required:

a) What is the probability that a new SME is successful?


1
Expert's answer
2022-05-17T15:24:13-0400

Let "A" denote the event "SME is started by graduate".

Let "B" denote the event "SME is started by non-graduate".

Let "S" denote the event "SME is successful".


"P(S)=P(A)P(S|A)+P(B)P(S|B)"

"=0.55(0.7)+0.45(0.1)=0.43"